Curious as to how to get started with the following types of problems 9/k+3 =3/5. How do I find the value of K?
Answer by Maths68(1474)

9/k+3 =3/5
9/k+3-3 =3/5-3
9/k=3/5-3
9/k=(3-15)/5
9/k=(-12)/5
9/k=(-12)/5
k/9=-5/12
k=-(5*9)/12
k=-45/12
k=-15/4

Check
=====
9/k+3 =3/5
9/(-15/4)+=3/5
9*4/(-15)+3=3/5
-36/15+3=3/5
-12/5+3=3/5
(-12+15)/5=3/5
3/5=3/5